A Contrived L^{A}TEX Example

                        Mark Senn <mark@purdue.edu>

September 30, 2004

Abstract

This document demonstrates L^{A}TEX.

1 Mostly Math

You can refer to things before they appear using \ref{label}. Distance is defined in equation 1.

See [2] and then [3] or [1] for L^{A}TEX information. Don’t do “d = \sqrt{(x_{2} − x_{1})^{2} + (y_{2} − y_{1})^{2}}.” In-

stead, if you have that much to say, use display math like

d = \sqrt{(x_{2} − x_{1})^{2} + (y_{2} − y_{1})^{2}}

Or, if you want to number it so you can refer to it somewhere else do

d = \sqrt{(x_{2} − x_{1})^{2} + (y_{2} − y_{1})^{2}}  (1)

The sum of the first ten integers, i.e., 1–10, is 55. The sum of the integers from 1 to n can be done “by hand.”

∑_{1}^{n} = 1 + 2 + ⋯ + n

= \frac{n(n + 1)}{2}  (2)

2 Figures

Figures automatically “float” to where L^{A}TEX thinks they’ll look good.

Sometimes input and output aren’t in same order.

Figure 1: Purdue “signature” logo

3 Tables

Tables also “float.” Use “[h]” to disable floating and put “here.”

Phone List

Name Phone
Aaron 1111
Zach 9999

Table 1: Phone List

References

[1] Helmut Kopka and Patrick W. Daly. Guide to L^{A}TEX. Addison-Wesley, Boston, Massachusetts, fourth edition, 2004.

[2] Leslie Lamport. L^{A}TEX: A Document Preparation System. Addison-Wesley, Reading, Massachusetts, second edition, 1995.

[3] Frank Mittelbach and Michel Goossens. The L^{A}TEX Companion. Addison-Wesley, Boston, Massachusetts, second edition, 2004.